40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 10 16:52:04 2020, gautam, Update, BHD, Workable setup prepared 
    Reply  Tue Aug 11 23:20:29 2020, gautam, Update, BHD, Some first tests with air BHD setup PRMI_RFlock.pdfPRMI_RFlock_fullscale.pdf
       Reply  Fri Aug 21 15:27:00 2020, gautam, Update, BHD, Better BHD mode-matching IMG_7548.JPGBHD_MM.pngPRMI_DCPDs.pdf
          Reply  Wed Aug 26 00:52:55 2020, gautam, Update, BHD, BHD activities 
Message ID: 15540     Entry time: Wed Aug 26 00:52:55 2020     In reply to: 15535
Author: gautam 
Type: Update 
Category: BHD 
Subject: BHD activities 

Listing some talking points from the last week of activity here.

  1. LO delivery fiber cable may be damaged.
    • The throughput itself doesn't suggest any problems, I get almost all the light I put in out the other end.
    • However, even when I slightly move the fiber, I see huge amplitude fluctuations in the DCPD readouts. This shouldn't be the case, particularly if the light is well matched to one of the special axes of the PM fiber. I checked with a PBS at the output that this is indeed the case, so something else must be funky?
    • In any case, I don't think it's a great idea to use this 70m long fiber for bringing the light from the PSL table to the adjacent AP table. Chub has ordered a 10m patch cable.
    • I was a bit too hasty this morning, thinking we had a patch cable in hand, and so I removed the fiber from the AP table. So right now, the LO beam doesn't make it to the BHD setup. Depending on the lead time for the new patch cable, I may or may not resurrect this old setup.
    • I have also located some foam and rigid plastic tubing which I think will help in isolating the fiber from environmental length(phase) modulation due to acoustic pickup.
  2. BHD commissioning activities
    • Basically, I've been trying to use the Single Bounce ITM reflection/ Michelson / PRMI with carrier locked to get some intuition about the BHD setup. These states are easily prepared, and much easier to understand than the full IFO for these first attempts.
    • One concern I have is the angular stability (or lack thereof). When the PRMI is locked, the DC light level on each DCPD fluctuates between ~0 (which is what it should be), up to ~30 cts (~85uW).
    • Using the empirically determined attenuation factor between the DCPDs and the dark port of the beamsplitter, I estimate the power can be as high as 20mW. This is a huge number, considering the input to the interferometer is ~800mW. I assume that all the light is at the carrier frequency, since the PRC should reject all the sideband light in this configuration. In any case, the total amount of sideband light is ~20mW, and the carrier stays resonant in the PRC even when there are these large ASDC excursions, so I think it's a reasonable assumption that the light is at the carrier frequency. Moreover, looking at the camera, one can see a clear TEM10/01 profile, indicative of imperfect destructive interference at the beamsplitter due to beam axis misalignment.
    • The effect of such excursions on the BHD readout hasn't yet been quantified (by me at least), but I think it may be hampering my attempts to dither the homodyne phase to estimate the LO phase noise.
  3. High voltage coil driver project - see thread for updates.
  4. Trek HV driver has arrived.
    • I haven't opened the box yet, but basically, what this means is that I can dither the mirror intended for homodyne phase control in a reasonable way.
    • Previously, I was using the OMC HV driver to drive the PZTs - but this dither signal path has a 2kHz high pass filter (since the OMC length dither is a kHz dither). I didn't want to futz around with the electronics, particularly since the unit was verified to be working.
    • So the plan now would be to drive the input of the Trek with a DAC output (an appropriate AI chassis has been prepared to interface with the CDS system).
    • Hopefully, there's enough DAC dynamic range to dither the PZT and also do the homodyne phase locking using a single channel. Else, we'd need to use two channels and install a summing amplifier.
    • We definitely need more high-voltage amplifiers/supplies in the lab:
      • Any Thorlabs HV drivers we can recover? 
      • Eventually, we will need HV for coil drivers, OMC PZTs, steering PZTs, homodyne phase control PZT. 
  5. PMC bases have arrived.
    • Joe Benson from the machine shop informed me today afternoon that the bases were ready for pickup.
    • We have 3 bases in hand now. The finish isn't the greatest in the world, but I think it'll work. You can see some photos here.
    • I will hold off on putting this together while I work on the basic airBHD commissioning tests. We can install the PMCs later.
  6. AS port WFS project
    • We now have in hand almost all the components for stuffing the ISC whitening and LSC demod boards.
    • Rich, Chub, Luis and I had a call on Monday. The advise from Rich/Luis was:
      • Choose an inductance that has Z~100 ohms at the frequency of interest, for the resonant transimpedance part.
      • Choose a capacitance that gives the appropriate resonant frequency.
      • Don't stuff more notches than you need - start with just a 2f notch (so 110 MHz for us), and make sure to place the highest frequency notch closest to the photodiode.
      • Rich also suggested looking at the optical signal with a non-optimized head, get an idea of what the field content is, and then tune the circuit as necessary. There are obviously going to be many issues that only become apparent once we do such a test.
    • The aLIGO modulation frequencies are only 20% different from the 40m modulation frequencies. So I thought it is best if for our first pass, we stick to the inductance values used in the aLIGO circuits (same footprint, known part etc etc). Then, we will change the capacitance so that we have a tuning range that is centered our modulation frequencies.
    • The parts have been ordered.
  7. ISS project
    • Half of the LO light on the BHD breadboard is diverted for the purpose of sensing the LO intensity noise, for eventual stabilization. Right now, it is just getting dumped.
    • A PD head has been located. It has a minimalist 1kohm transimpedance amplifier circuit integrated into the head.
    • Our AOM driver has an input range of 0-1V DC. We want to map the servo output of +/-10V DC (or +/-4V DC if we use an SR560 based servo for a first pass) to this range.
    • I wanted to do this for once in a non-hacky way so I drew up a circuit that I think will serve the purpose. It has been fabricated and will be tested on the bench in a couple of days.
    • Once I get a feel for what the signal content is, I will also draw up a interface board to the PD head that (i) supplies the reverse bias voltage and +/-15 V DC to the PD head and (ii) applies some appropriate HPF action and provides a DC monitor as well.
  8. Summary pages are dead.
  9. General lab cleanup
    • I moved all the PPE from the foyer area into the designated cabinets along the east arm.
    • Did some basic cleanup of the lab in preparation for crane inspection. Walkways are clear.
    • I de-cluttered the office area a bit, but today I received ~10 packages from Digikey/FrontPanelExpress etc. So, in fact, it got even more cluttered. Entropy will go down once we ship these off to screaming circuits for stuffing the PCBs.
ELOG V3.1.3-